Prove that the following series is convergent for all 𝑟 ∈ ℝ

Σ (1 + 1/2 + ... + 1/n) (sin (nr) / n).



1
Expert's answer
2022-01-17T16:56:46-0500

let:

"a_n=\frac{Σ ( 1/n)}{ n}|sin(nr)|"


"b_n=\frac{Σ ( 1/n)}{ n}"


"|sin (nr)|\le 1" , then:

"a_n\le b_n"


also:

"b_n\le \sum (1/n^2)"


So, since "\sum (1/n^2)" converges, "b_n=\frac{Σ ( 1/n)}{ n}" converges as well

so, "\sum b_n" converges

then, "\sum a_n" converges


so, series "\sum (1+1/2+...+1/n)\frac{sin(nr)}{n}" converges